y^2logy=xpylogx+x^2p^2 by using logy=v,logx=u find G.S,S.S
1
Expert's answer
2020-10-26T19:46:19-0400

Given equation is "y^2logy=xpylogx+x^2p^2"

Putting, "log y = v, logx = u"

then,

"\frac{1}{y}dy = dv, \frac{1}{x}dx = du" "\implies \frac{dy}{dx} = \frac{y}{x}\frac{dv}{du}"


Putting the values, then equation will be,

"y^2(\frac{dv}{du})^2+xyu\frac{y}{x}\frac{dv}{du} - y^2v = 0"


"(\frac{dv}{du})^2+u\frac{dv}{du} - v = 0"


Let, "P = \frac{dv}{du}" then "P^2 +uP-v = 0"

Differentiating equation, we get,

"2P\frac{dP}{du}+P+u\frac{dP}{du}-P = 0 \implies \frac{dP}{du}(2P+u) = 0"


Then, "\frac{dP}{du} = 0 \implies P = k"

"\frac{dv}{du} = k \implies v = ku" (1)



Also, "2P+u = 0 \implies 2dv = -udu"

"2v = -\frac{1}{2}u^2+C" (2)


From (1), "e^y = ke^x \implies y =Kx" (3)


From (2), "2e^y = -\frac{1}{2}e^{2x} + C" (4)



Equation(3) and (4) are the required solutions.